# Generative Inverse Design of Metamaterials with Functional Responses by Interpretable Learning

### v4

@@ 4.3 Diversity versus Feasibility of Generated Designs
The MCMC in RIGID can sample new designs in the entire design space based on the estimated likelihood, which is not limited to local optimality and has the potential to discover multiple very different satisfying solutions (if any) in one shot. This is clearly demonstrated in the SupSin example where there are multiple disconnected feasible regions in the design space and the likelihood estimated by RIGID successfully coincides with these feasible regions (Fig. 7A). The ability to generate multiple satisfying solutions allows us to consider other figures of merit, such as manufacturability and cost, in addition to the main objective of satisfying the target response. We can also control the likelihood threshold to adjust the diversity of generated designs.
For all four problems demonstrated in this work, when the sampling threshold is set to 0.8, RIGID achieves a satisfaction rate of 100%. Although GA can also lead to a high target satisfaction rate for relatively simple inverse design problems (e.g., the optical metasurface design problem in Sec. 3.2), the generated solutions lack diversity. In contrast, RIGID provides a unique way to control the trade-off between the diversity and the feasibility of generated designs by tuning the sampling threshold. It is important to note that, without considering diversity, a high target satisfaction rate becomes trivial, as we can simply apply small perturbations to one feasible solution to generate an infinite number of similar solutions that have similar responses and hence are also likely to meet the target. In many inverse design problems where solutions are non-unique, finding one solution can be simple, but the real challenge lies in discovering a broad range of feasible solutions—a problem where RIGID will outperform optimization-based methods as demonstrated by our results.
@@ 4.4 Discrete Design Variables
The use of a random forest in RIGID allows easy consideration of discrete design variables. For example, our optical metasurface design problem has the geometry type and the material choice for each layer as categorical design variables. In contrast, these discrete variables will pose challenges for other inverse design methods based on some machine learning models (e.g., conditional generative models) or gradient-based optimization. On the other hand, the simplicity of the tree-based model does not hurt its capability of solving complex problems, as a decision tree can represent arbitrarily complex decision boundaries.
@@ Acknowledgement
Appendix S1 Existing Works on Deep Learning-Based Iteration-Free, Single-Shot Inverse Design
@@
Figure S7: Designs and corresponding absorbance spectra of generated optical metasurface designs for design targets \@slowromancapii@-\@slowromancapiv@ (marked as gray-shaded wavelength regions). Each row shows five designs with the highest likelihood of satisfying a target. All the layer thicknesses ($h_{l},l=1,2,3$) have a unit of nm.
Figure S8: Designs and corresponding absorbance spectra of generated optical metasurface designs for design targets \@slowromancapv@-\@slowromancapvii@ (marked as gray-shaded wavelength regions). Each row shows five designs with the highest likelihood of satisfying a target. All the layer thicknesses ($h_{l},l=1,2,3$) have a unit of nm.
Figure S9: Designs and corresponding absorbance spectra of generated optical metasurface designs for design targets \@slowromancapviii@-\@slowromancapx@ (marked as gray-shaded wavelength regions). Each row shows five designs with the highest likelihood of satisfying a target. All the layer thicknesses ($h_{l},l=1,2,3$) have a unit of nm.
@@
Figure S11: Examples of generated designs that satisfy the targets and belong to the rectangular ring and H-shaped geometry types. All the layer thicknesses ($h_{l},l=1,2,3$) have a unit of nm.

### v6

@@ 4.3 Diversity versus Feasibility of Generated Designs
The MCMC in RIGID can sample new designs in the entire design space based on the estimated likelihood, which is not limited to local optimality and has the potential to discover multiple very different satisfying solutions (if any) in one shot. This is clearly demonstrated in the SupSin example where there are multiple disconnected feasible regions in the design space and the likelihood estimated by RIGID successfully coincides with these feasible regions (Fig. 7A). The ability to generate multiple satisfying solutions allows us to consider other figures of merit, such as manufacturability, compatibility, and cost, in addition to the main objective of satisfying the target response. Besides enabling flexibility of design choice, generating a comprehensive solution set can also bring more insights into the structure-property relation associated with the target. For example, after obtaining a diverse set of designs satisfying the same target, we can investigate the differences among these designs to identify critical features that lead to the satisfaction of the target.
For all four problems demonstrated in this work, when the sampling threshold is set to 0.8, RIGID achieves a satisfaction rate of 100%. Although GA can also lead to a high target satisfaction rate for relatively simple inverse design problems (e.g., the optical metasurface design problem in Sec. 3.2), the generated solutions lack diversity. In contrast, RIGID provides a unique way to control the trade-off between the diversity and the feasibility of generated designs by tuning the sampling threshold. It is important to note that, without considering diversity, a high target satisfaction rate becomes trivial, as we can simply apply sufficiently-small perturbations to one feasible solution to generate an infinite number of similar solutions that have similar responses and hence are also likely to meet the target. In many inverse design problems where solutions are non-unique, finding one solution can be simple, but the real challenge lies in discovering feasible solutions that cover a broad range of the design space—a problem where RIGID will outperform optimization-based methods as demonstrated by our results.
@@ 4.4 Discrete Design Variables
The use of a random forest in RIGID allows easy consideration of discrete design variables. For example, our optical metasurface design problem has the geometry type and the material choice for each layer as categorical design variables. In contrast, these discrete variables will pose challenges for other inverse design methods based on some machine learning models (e.g., conditional generative models) or gradient-based optimization.
@@ 6 Experimental Section
Configurations for GA
The initial population for GA is generated by randomly sampling solutions within the design variable bounds. At each generation, we select individuals to reproduce for a new generation by randomly picking three individuals from the population, selecting the best one, and repeating this process 300 times. We set the probability for crossover and mutation to be 0.5 and 0.2, respectively. Once an individual is mutating, there is a 5% chance for each design variable in this individual to mutate. For a continuous design variable, we define the mutation operation as adding Gaussian noise with a mean of 0 and a standard deviation equal to 1% of the variable’s range. For a categorical design variable, mutation is defined as randomly reassigning the category of the variable. We ran GA for 10 generations, each with a population of 300. Finally, we select a set of solutions with the highest objective values among all the individuals created during GA, and use these selected solutions as generated designs.
@@ Acknowledgement
Appendix S1 Existing Approaches on Deep Learning-Based Iteration-Free, Single-Shot Inverse Design
@@
Figure S7: Designs and corresponding absorbance spectra of generated optical metasurface designs for design targets \@slowromancapii@-\@slowromancapiv@ (marked as blue-shaded wavelength regions). Each row shows five designs with the highest likelihood of satisfying a target. All the layer thicknesses ($h_{l},l=1,2,3$) have a unit of nm.
Figure S8: Designs and corresponding absorbance spectra of generated optical metasurface designs for design targets \@slowromancapv@-\@slowromancapvii@ (marked as blue-shaded wavelength regions). Each row shows five designs with the highest likelihood of satisfying a target. All the layer thicknesses ($h_{l},l=1,2,3$) have a unit of nm.
Figure S9: Designs and corresponding absorbance spectra of generated optical metasurface designs for design targets \@slowromancapviii@-\@slowromancapx@ (marked as blue-shaded wavelength regions). Each row shows five designs with the highest likelihood of satisfying a target. All the layer thicknesses ($h_{l},l=1,2,3$) have a unit of nm.
@@
Figure S11: Examples of generated designs that satisfy the targets (marked as blue-shaded wavelength regions) and belong to the rectangular ring and H-shaped geometry types. All the layer thicknesses ($h_{l},l=1,2,3$) have a unit of nm.